# Improved finite-size key rates for discrete-modulated continuous variable quantum key distribution under coherent attacks

### v1

@@ IV.4 Optimizing the Min-Tradeoff Function
In this section we describe how we can transform the optimization problem given by eq. (92) to obtain an affine min-tradeoff function. Note that this optimisation involves an infinite-dimensional space on Bob’s side. In order to solve this inconvenience for the numerics, we will introduce a cutoff in the number of photons Bob receives, $N_{c}\in\mathbb{N}$, and solve the resulting optimisation for increasing values of $N_{c}$. As the obtained results numerically converge, that is, at some point hardly change when increasing $N_{c}$, we will make a cutoff assumption and assume that the solution obtained for a large enough $N_{c}$ is almost equal to the one obtained using an unbounded $N_{c}$. This convergence is a reasonable assumption, provided the separability of the underlying Hilbert space, as well as the exponential attenuation at the optical fiber—which ensures that only low photon number states will be neatly populated.
@@
The principle of this technique is reformulating the relative entropy as a convex cone [57], and using an efficient implementation [58] of the recently discovered Skajaa and Ye’s algorithm [43, 59] to optimize over it. While this allows for the solution of a wide class of optimization problems involving the relative entropy, it is not yet enough to solve (92). The difficulty is that this problem is not strictly feasible, as the matrix $\mathcal{G}(\omega_{AB})$ will necessarily have null eigenvalues, and strict feasibility is necessary for the algorithm to work reliably. We emphasize that this is not a limitation of this particular method, but a generic feature of convex optimization algorithms [60, 61].
@@
The first inequality is due to the fact that $r$ is a convex, differentiable function over the convex set $\mathcal{D}(\mathcal{H}_{AB})$, hence it can be lower bounded by its first order Taylor expansion at $\tilde{\omega}$ (see e.g. [62, p.69]), and the last inequality is due to the fact that $\omega_{\tilde{p}}^{*}\in\Sigma^{\prime}(\tilde{p})$ is the primal optimal. On the other hand, we calculate the matrix gradient of the objective function
@@
for any $\tilde{\omega}$ [63], so we can effectively remove such term in (99). Now, for any $\tilde{\omega}$ and $\tilde{p}$, the optimisation problem in eq. (99) is an instance of semidefinite program (SDP) in standard form, explicitly given by
@@ V Numerical implementation and results
Here, the different integration intervals represent the fragments of the phase space corresponding to the modulation expressed in (17), and $\phi_{x}\in\{i^{x}\alpha\}_{x=0}^{3}$ are the coherent state amplitudes used by Alice with $\alpha\in\mathbb{R}$. The region operators for the constraints $\{\tilde{R}_{B}^{z}\}_{z}$ are given by (50), such that they require an integration in the same, corresponding intervals. As it will be the case for the states, the numerical analysis requires us to project the operators onto a Fock basis bounded by a cutoff $N_{c}$, where values $N_{c}\geq 10$ are typically quoted [28, 61]. We note that the cutoff represents a reasonable assumption, since the states are bounded in their energies according to the attenuation at the channel (so that high photon states are unoccupied), and that the underlying Hilbert space is separable—such that any infinite dimensional state can be well approximated by an appropriate finite basis. Using then the inner product [65]
@@ VI Discussion
This improvement in the finite-size regime can be partly attributed to the better performance of the new numerical method [44], as well as the use of the GEAT in combination with [42] instead of the original EAT. In [37], the EAT has to be combined with a virtual tomography of Alice’s subsystem in order to take into account the fact that the protocol is prepare-and-measurement based, whereas the approach of [42] allows for a direct application of the GEAT to said protocol. Since such virtual tomography requires not only extra steps in the security proof but also an expenditure in the key rate via statistical estimators, we thus achieve a simplified security framework as well as an increase in the finite key rate. Such advancement, however, comes with one caveat. Namely, the method of [42] requires that Eve is only allowed to be in possession of one state at a time, which can be enforced by Alice waiting for Bob’s confirmation of having received a state before sending the next state. Since this has to be performed in every round, this might slow down the protocol. How big this effect is depends on the actual implementation. The method of virtual tomography [37], which could also be improved by using the numerical approach of [44], does not make any assumptions on the number of states Eve might hold at a time, thus not requiring a confirmation by Bob before Alice sends the next state. This shows that the two methods are somewhat complimentary. Very recently, it has been noted (see [72, Remark 8]) that it might be possible to lift the condition of Eve only having one subsystem at a time by combining the GEAT approach, which we use in this work, with the approach of [72, 73].
@@
Another point to consider is the photon number cutoff. Whereas our security proof provided here already allows for infinite dimensional Hilbert spaces, our numerics still assume a maximal photon number of $N_{c}$. There have been several results which overcome this assumption in the scenario of collective attacks [29, 32, 33] by splitting (92) into an optimisation over states on a finite Hilbert space and a correction term—using an energy test, which can be seen as an additional part of the parameter estimation, one can bound the probability of being outside the cutoff space with high probability. The bound on the probability then goes into the higher order term, as well as in the constraints of the finite-size optimisation. Considering the applicability of said model in the case of coherent attacks, we have realised that combining such an approach with the GEAT poses additional challenges in the setting of finite block sizes. When applying the GEAT, the outputs of the hypothetical energy tests have to be included into the alphabet $\tilde{\mathcal{C}}$, becoming a set of variables in the min-tradeoff function. As the dependence on these probabilities is not affine, the min-tradeoff function has to be linearised by taking a first order Taylor expansion. This is possible in principle, but did not provide us with positive key rates due to emerging numerical difficulties. We therefore leave the lifting of the photon number cutoff assumption for future work. In particular, the recently announced Rényi version of both the EAT and GEAT [72] looks promising in that respect, as it does not require the definition of an affine min-tradeoff function.
@@ VII Acknowledgments
C.P.G thanks Marco Túlio Quintino for fruitful indications about numerical precision. S.B. thanks Omar Fawzi, Min-Hsiu Hsieh, Florian Kanitschar, Bill Munro, Mizanur Rahaman, Gelo Noel Tabia, Toshihiko Sasaki and Shin-Ichiro Yamano for insightful discussions. This work was supported by the ERC (AdG CERQUTE, grant agreement No. 834266), the AXA Chair in Quantum Information Science, Gobierno de España (Severo Ochoa CEX2019-000910-S, NextGen Quantum Communications and FUNQIP), Fundació Cellex, Fundació Mir-Puig, the EU (QSNP and Quantera Veriqtas), the Generalitat de Catalunya (CERCA program and the postdoctoral fellowship programme Beatriu de Pinós), European Union’s Horizon 2020 research and innovation programme under grant agreement No. 801370 (2019 BP 00097) within the Marie Sklodowska-Curie Programme. The research of M.A. was supported by the European Union–Next Generation UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León.

### v2

@@ IV.4 Optimizing the Min-Tradeoff Function
In this section we describe how we can transform the optimization problem given by eq. (92) to obtain an affine min-tradeoff function. Note that this optimisation involves an infinite-dimensional space on Bob’s side. In order to solve this inconvenience for the numerics, we will introduce a cutoff in the number of photons, $N_{c}\in\mathbb{N}$, and solve the resulting optimisation for increasing values of $N_{c}$. As the obtained results numerically converge, that is, at some point hardly change when increasing $N_{c}$, we will make a cutoff assumption and assume that the solution obtained for a large enough $N_{c}$ is almost equal to the one obtained using an unbounded $N_{c}$. This convergence is a reasonable assumption, provided the separability of the underlying Hilbert space.
@@
The principle of this technique is reformulating the relative entropy as a convex cone [58], and using an efficient implementation [59] of the recently discovered Skajaa and Ye’s algorithm [43, 60] to optimize over it. While this allows for the solution of a wide class of optimization problems involving the relative entropy, it is not yet enough to solve (92). The difficulty is that this problem is not strictly feasible, as the matrix $\mathcal{G}(\omega_{AB})$ will necessarily have null eigenvalues, and strict feasibility is necessary for the algorithm to work reliably. We emphasize that this is not a limitation of this particular method, but a generic feature of convex optimization algorithms [61, 62]. [^3] We also note that $\omega$ itself does not have null eigenvalues according to the nature of our numerical method. Namely, the numerical solver starts with an initial, full rank state and performs the optimization according to a barrier function, whose design effectively prevents the states from having non-positive eigenvalues.
@@
The first inequality is due to the fact that $r$ is a convex, differentiable function over the convex set $\mathcal{D}(\mathcal{H}_{AB})$, hence it can be lower bounded by its first order Taylor expansion at $\tilde{\omega}$ (see e.g. [63, p.69]), and the last inequality is due to the fact that $\omega_{\tilde{p}}^{*}\in\Sigma^{\prime}(\tilde{p})$ is the primal optimal. On the other hand, we calculate the matrix gradient of the objective function
@@
for any $\tilde{\omega}$ [64], so we can effectively remove such term in (99). Now, for any $\tilde{\omega}$ and $\tilde{p}$, the optimisation problem in eq. (99) is an instance of semidefinite program in standard form, explicitly given by
@@
IV.5 Assumptions for security
Before moving to the computation of key rates, we would like to comment on the assumptions used in the derivation of our security proof. First of all, the considered protocol is formulated in a fully device-dependent manner and, therefore, assumes trusted state preparations by Alice and measurements by Bob, as in the vast majority of previous CVQKD security proofs. In our case, we also use two additional assumptions: (i) a bounded-energy assumption, where Eve’s attack uses bounded, but arbitrary, energy; and (ii) a cut-off assumption, saying that the asymptotic trade-off computed for large enough number of photons, $N_{c}$, is very close to the value for arbitrary $N_{c}$. This is not to be confused with the standard formulation of the cut-off assumption, usually presented as (i’) Eve’s attack uses states only with a bounded number of photons $N_{c}$. In particular our security proof under assumptions (i) and (ii) also implies security under assumption (i’), but the opposite is not necessarily true.
@@ V Numerical implementation and results
Here, the different integration intervals represent the fragments of the phase space corresponding to the modulation expressed in (17), and $\phi_{x}\in\{i^{x}\alpha\}_{x=0}^{3}$ are the coherent state amplitudes used by Alice with $\alpha\in\mathbb{R}$. The region operators for the constraints $\{\tilde{R}_{B}^{z}\}_{z}$ are given by (50), such that they require an integration in the same, corresponding intervals. As it will be the case for the states, the numerical analysis requires us to project the operators onto a Fock basis bounded by a cutoff $N_{c}$, where values $N_{c}\geq 10$ are typically quoted [28, 62]. It is observed that the obtained key rates numerically converge when increasing the cutoff, which is expected since the states are bounded in their energies according to the attenuation at the channel (so that high photon states are unoccupied), and that the underlying Hilbert space is separable—such that any infinite dimensional state can be well approximated by an appropriate finite basis. Using then the inner product [66]
@@ V.1 Comparison to other works
V.2 Comparison to practical implementations
The obtained improvements in terms of block sizes, however, come in principle with a caveat, namely, the method of GEAT [42] used in the security proof requires that Eve is only allowed to be in possession of one state at a time. Since this condition must be satisfied in every round, the repetition rate of the protocol clearly slows down, to enforce that there is only one signal in the channel at a time. For instance, the speed of light in an optical fiber is $\tilde{c}\approx 2c/3$, for $c$ the one in vacuum, which for distances of the order of 20 km implies a repetition rate of the order of 10 Khz. This is however not a particularly relevant limitation for current CVQKD implementations. In fact, the overall execution time in practical CVQKD implementations is fundamentally limited by the postprocessing of the signals – such that, as in the case of a recent work [72], it takes of the order of 3 minutes to perform the digital signal processing for every frame of $10^{6}$ rounds. This indicates, provided a full post-processing of all the frames in parallel, a repetition rate of $\nu\approx 5.6\times 10^{3}$ Hz, slower than the one imposed by the GEAT requirements. Thus, using the GEAT does not necessarily represent a limitation provided current technological standards, as long as Alice and Bob perform a postprocessing of the measured signals while also carrying the preparation and measurement of the remaining ones.
Figure 6 provides an estimation of the secret key rates in terms of bits per pulse, where the repetition rate $\nu$ is fixed according to the maximal value achieved due to the postprocessing time of [72], and the repetition rate induced by the sequential structure of the GEAT – namely, $\tilde{\nu}=\tilde{c}/L$ where $L$ is the distance between Alice and Bob. The latter represents the ultimate rate provided by our analysis when the GEAT sequential requirement is the main limitation, while the former gives the expected rates when applying our formalism to state-of-the art implementations.
$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{5}$051015202530Refer to caption
Secret Key Rate (bits/sec)
Distance (km)$n=1\mathrm{e}9$, $\tilde{\nu}$$n=1\mathrm{e}9$, $\nu$$n=5\mathrm{e}8$, $\tilde{\nu}$$n=5\mathrm{e}8$, $\nu$
Figure 6: Finite secret keys as bits per unit time according to different block sizes $n$ and repetition rates $\nu$ according to the postprocessing time [72], and $\tilde{\nu}=\tilde{c}/L$ the sequential constraint imposed by the GEAT, with $\tilde{c}$ the speed of light in the fiber and $L$ the distance. Any other parameters were taken to be the same as in Figure 3.
@@ VI Discussion
This improvement in the finite-size regime can be partly attributed to the better performance of the new numerical method [44], as well as the use of the GEAT in combination with [42] instead of the original EAT. In [37], the EAT has to be combined with a virtual tomography of Alice’s subsystem in order to take into account the fact that the protocol is prepare-and-measurement based, whereas the approach of [42] allows for a direct application of the GEAT to said protocol. Since such virtual tomography requires not only extra steps in the security proof but also an expenditure in the key rate via statistical estimators, we thus achieve a simplified security framework as well as an increase in the finite key rate.
This enhanced performance comes with the structural limitations of the GEAT, where the repetition rate of the protocol is constrained by the requirement that Eve holds only one register at every step. As we observed from [72], this is not critical for current experimental platforms, as long as the measurements are postprocessed during the preparation and measurement of the signals. However, future implementations may provide higher key rates by means of new techniques (e.g. code parallelization and precompilation, GPU-optimized performance), although said techniques typically come at a sharp cost and higher energy consumption. In this context, it also has been noted (see the discussion on page 3, and Corollary 6.1 of [74]) that it might be possible to lift the condition of Eve only having one subsystem at a time by combining the GEAT approach, which we use in this work, with the approach of [75]. On similar grounds, the method of virtual tomography [37], which could also be improved by using the numerical approach of [44], does not make any assumptions on the number of states Eve might hold at a time.
@@
Another point to consider is the photon number cutoff. Whereas our security proof provided here already allows for infinite dimensional Hilbert spaces, our numerical computation of tradeoff functions, or asymptotic key rates, fixes a maximal photon number of $N_{c}$, which is however increased until observing numerical convergence. There have been several results which overcome this assumption in the scenario of collective attacks [29, 32, 33] by splitting (92) into an optimisation over states on a finite Hilbert space and a correction term—using an energy test, which can be seen as an additional part of the parameter estimation, one can bound the probability of being outside the cutoff space with high probability. The bound on the probability then goes into the higher order term, as well as in the constraints of the finite-size optimisation. Considering the applicability of said model in the case of coherent attacks, we have realised that combining such an approach with the GEAT poses additional challenges in the setting of finite block sizes. When applying the GEAT, the outputs of the hypothetical energy tests have to be included into the alphabet $\tilde{\mathcal{C}}$, becoming a set of variables in the min-tradeoff function. As the dependence on these probabilities is not affine, the min-tradeoff function has to be linearised by taking a first order Taylor expansion. This is possible in principle, but did not provide us with positive key rates due to emerging numerical difficulties. We therefore leave the lifting of the photon number cutoff assumption for future work. In particular, the recently announced Rényi version of both the EAT and GEAT [74] looks promising in that respect, as it does not require the definition of an affine min-tradeoff function. Another possible avenue of reserach that could yield even better rates would be to use the techniques applied in [64, 76] to optimise the min-tradeoff function.
@@ VII Acknowledgments
C.P.G thanks Marco Túlio Quintino for fruitful indications about numerical precision, and Yoann Piétri for suggestions about experimental aspects of CVQKD. We further thank Omar Fawzi, Min-Hsiu Hsieh, Lars Kamin, Florian Kanitschar, Bill Munro, Mizanur Rahaman, Gelo Noel Tabia, Ernest Tan, Toshihiko Sasaki and Shin-Ichiro Yamano for insightful discussions. This work was supported by the ERC (AdG CERQUTE, grant agreement No. 834266), the AXA Chair in Quantum Information Science, Gobierno de España (Severo Ochoa CEX2019-000910-S, NextGen Quantum Communications and FUNQIP), Fundació Cellex, Fundació Mir-Puig, the EU (QSNP and Quantera Veriqtas), the Generalitat de Catalunya (CERCA program and the postdoctoral fellowship programme Beatriu de Pinós), European Union’s Horizon 2020 research and innovation programme under grant agreement No. 801370 (2019 BP 00097) within the Marie Sklodowska-Curie Programme. The research of M.A. was supported by the European Union–Next Generation UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León, and by the Spanish Agencia Estatal de Investigación, Grant No. RYC2023-044074-I.